# SOP-Bench: Complex Industrial SOPs for Evaluating LLM Agents

### v1

@@ 5 Evaluating Agents with SOP-Bench
We evaluate FC-Agent and ReAct-Agent on SOP-Bench using three inputs: (1) Task, (2) SOP (instructions), and (3) ToolSpecs. The outputs include (1) intermediate tool outputs, (2) final response, and (3) trace of tool use and reasoning. We define: $n$ (total number of tasks per SOP), $n_{c}$ (number of tasks per SOP marked complete by the agent), and $n_{a}$ (number of correctly completed tasks per SOP). Evaluation metrics are: (1) Execution Completion Rate ($ECR$): proportion of tasks the agent marked as complete, $ECR=\frac{n_{c}}{n}$; (2) Conditional Task Success Rate ($C$-$TSR$): fraction of completed tasks that match the ground truth, $C$-$TSR=\frac{n_{a}}{n_{c}}$; and (3) Task Success Rate ($TSR$): overall accuracy, $TSR=\frac{n_{a}}{n_{t}}$. While $ECR$ and $TSR$ serve as core metrics, certain tasks demand domain-specific evaluation. For example, content moderation emphasizes false positive/negative rates, while patient intake prioritizes process adherence and compliance with privacy standards. Though this work reports only $ECR$ and $TSR$, future research should incorporate tailored metrics to better assess agent performance across industry verticals.
Table 4: Agent Evaluation Results on SOP-Bench: ECR = Execution Completion Rate; TSR = Task Success Rate; C-TSR = Conditional Task Success Rate. All percentage values are rounded up to the nearest integer.
SOP Topic $n$ FC-Agent React-Agent
ECR C-TSR TSR ECR C-TSR TSR
Content Flagging 226 0% 0% 0% 90% 11% 10%
Customer Service 208 100% 14% 14% 88% 4% 3%
Dangerous Goods 327 68% 87% 59% 83% 96% 80%
Aircraft Inspection 150 100% 1% 1% 97% 89% 87%
Email Intent 122 78% 39% 31% 60% 100% 60%
Know Your Business 122 100% 19% 19% 92% 39% 36%
Patient Intake 90 33% 0% 0% 100% 0% 0%
Video Annotation 168 100% 49% 49% 70% 99% 69%
Video Classification 198 100% 52% 52% 54% 39% 21%
Warehouse Package Inspection 200 41% 32% 13% 54% 39% 21%
Weighted Average 70% 35% 27% 83% 61% 48%
Figure 3: Shows the relationship between human perceived complexity and task success rates (TSR) of the two agents. There is very low correlation, $\rho(\text{TSR}_{\text{FC}},\mathbb{C_{H}})=0.19$, $\rho(\text{TSR}_{\text{React}},\mathbb{C_{H}})=0.12$, both $p>0.10$, so statistically insignificant, thus highlighting the fact that even SOPs with low human perceived complexity could be complex for LLMs.
Table 4 shows the average  [^3] We have averaged percentages of agents’ success rates over 10 SOPs. With each SOP being different from the others we dont think the error distribution is identical, which violates the normality assumption. Hence, we are not reporting error bars of the average proportion estimates. performance metrics of both FC-Agent and React-Agent on all the tasks in SOP-Bench. On average, ReAct-agent performs better than FC-agent. Such analysis and conclusions are only possible when a benchmark like SOP-Bench is available.
6 Discussion
In explaining our preliminary findings/results, we consider several factors that could affect agent performance, such as (1) number of tools needed, (2) complexity of reasoning required, (3) presence of ambiguous or implicit SOP instructions, (4) need for long-term memory or state tracking, and (5) degree of procedural vs. declarative knowledge involved. For instance, React-Agent performs exceptionally well on tasks like Dangerous Goods Classification SOP and Email Intent SOP, which require nuanced reasoning and decision chaining, but fails completely on tasks like Customer Service and Patient Intake that demand precise form filling and tool invocation—highlighting its difficulty with structured, deterministic workflows. Conversely, the FC-Agent succeeds on tasks where deterministic tool use is sufficient, but struggles when the task requires deeper context inference or adaptive reasoning, as seen in Content Flagging or Know Your Business.
The content flagging SOP which resulted in lowest ECR and TSR, involves four API calls, each requiring five to six input parameters and a strict execution order. The agent must reliably invoke these tools in the correct sequence, accurately populate the input parameters, and retain intermediate outputs in memory for use in subsequent steps. However, analysis of the agent logs reveals frequent failures due to tool invocation errors and inconsistent function outputs. Despite aiming to follow the SOP closely, the agent often deviates—either by skipping steps or manually approximating values when a tool fails. In several instances, the agent resorts to hallucinating results; for example, when the calculate user trust score function fails, it fabricates a value between 0 and 100. These issues lead to lower accuracy (lower TSR).
The evaluation of tool-calling accuracy on two SOPs shows significant variability, highlighting both the challenges and potential of using LLMs for task automation. In the Content Flagging SOP, which uses a smaller set of four tools, the FC-agent successfully invoked all required tools in 66.9% of tasks. However, this includes cases with incorrect parameters, exposing a gap in parameter handling. Despite the high invocation rate, 79.3% of tool calls failed, with 60.6% of failures due to parameter issues like sequence mismatches or invalid argument types. These findings underscore the need for better parameter validation and error handling to enhance reliability.
The Video Classification SOP posed a more complex challenge, with 25 available tools but only 5 required per task. Despite this, the agent invoked at least one incorrect tool in every execution, leading to a 100% rate of unwanted tool calls. Although the agent eventually completed the tasks, frequent irrelevant tool use reduced efficiency and reliability. Overall, 74.8% of tool calls failed, with 50.6% due to incorrect parameters, highlighting the difficulty of parameter alignment in complex, multi-tool settings. As task complexity grows, accurate tool selection and parameter handling become increasingly critical for reliable execution.
Finally, the near-zero TSR for both agents in Patient Intake illustrate another critical challenge: tasks involving sensitive domains (e.g., healthcare) often require strict sequencing, nuanced input parsing, and adherence to implicit compliance policies. Current agents struggle to manage these interleaved constraints without explicit fine-tuning or specialized architectural support.
7 Conclusions and Next Steps
This work presents SOP-Bench, a comprehensive benchmark generation workflow and associated benchmarks that are designed to evaluate AI agents on real-world SOPs across diverse industrial domains. Our preliminary findings with two popular agent architectures validate the central thesis of this work: agent performance is highly context- and task-dependent, and robust benchmarking across realistic SOPs is essential to inform further progress and research.
In a nutshell, SOP-Bench represents a significant leap forward in benchmarking the performance and robustness of AI agents in enterprise automation settings. It enables rigorous evaluation of agents on realistic, industrial-grade SOPs that reflect real-world complexity including task inter-dependencies, procedural ambiguity, conditional logic, and error-handling requirements. Moving forward, there are several promising directions for extending this benchmark:
(1) Introduce multiple versions of the same SOP with varying levels of complexity and abstraction and systematically study how agent performance scales with task difficulty.
(2) Incorporate multiple LLMs in the SOP and data generation pipeline to reduce linguistic and structural bias, and to add richness to procedural variations.
(3) Evaluate additional agent architectures—including retrieval-augmented agents, multi-agent collaborative systems, and memory-augmented agents, to better understand trade-offs and strengths.
(4) Extend SOP-Bench to include multimodal instructions such as images and tables, which are common in enterprise documentation and workflows.
(5) Include hierarchical SOPs — with sections and sub-sections — that involve significant context switching back and forth, thereby adding substantial complexity to the SOPs.
We encourage the research community and practitioners to use SOP-Bench in two primary ways: (1) to evaluate and compare different agent designs for tasks that are structurally or semantically similar to those in their domain, and (2) to adopt our framework for generating task-specific SOP benchmarks when such examples are not readily available in SOP-Bench. In either case, systematic benchmarking should precede production deployment to ensure that the chosen architecture aligns with domain-specific requirements, robustness needs, and compliance constraints.
Broader impact
Positive Societal Impacts
SOP-Bench has the potential to significantly accelerate the deployment of safe and effective LLM agents in high-stakes industrial workflows, where reliability and consistency are paramount. Furthermore, the public release of SOP-Bench, along with the data generation framework, democratizes access to high-quality benchmarks for both academic and industry researchers. This can foster transparent development, reproducible evaluation, and community-driven progress toward safer AI systems for enterprise applications.
Negative Societal Impacts
There is a risk that widespread adoption of SOP-following agents may lead to over-reliance on automated decision-making in sensitive domains, such as healthcare or legal triage, without adequate human oversight. If not rigorously tested for edge cases and bias, LLM agents could propagate errors, reinforce institutional inequities, or introduce new failure modes in mission-critical systems.
We encourage responsible use of SOP-Bench, especially in regulated or high-risk domains, and advocate for human-in-the-loop solutions and evaluations in all downstream applications.
Disclaimer
IMPORTANT: The Standard Operating Procedures (SOPs) used in this benchmark are synthetically generated using Large Language Models for research and evaluation purposes only. These SOPs are designed to test AI agent capabilities and DO NOT represent actual operational procedures from any real organization. They should NOT be used or adapted for real-world implementation or deployment. The authors make no claims about the safety, completeness, or regulatory compliance of these synthetic SOPs. Similarly, all example data, including but not limited to patient records, user profiles, transaction logs, and any other structured or unstructured data used in this benchmark, are entirely synthetic and AI-generated. This data does not represent or correspond to any real individuals, organizations, or actual scenarios. Any resemblance to real persons, or organizations, or situations is purely coincidental. This synthetic data is provided solely for research purposes and should not be used as a basis for real-world decision-making or system implementation.
@@ Appendix A Technical Details on Agents
A.1 FC Agent
This is a prompt-driven, lightweight execution engine leveraging the native tool calling capability of LLM (e.g, Claude). It accepts an SOP text and a structured task input, dynamically generates prompts using a parameterized template engine, and maintains a conversation loop with Claude throughout the execution. During agent initialization, it registers external tools defined in an OpenAPI-style tools specifications JSON file. During execution, when the model triggers a tool_use event, the agent extracts the relevant tool name and parameters, executes the corresponding function via a domain-specific ToolsManager, and delivers the output to the model through a toolResult message. A built-in self-reflection mechanism automatically responds to non-critical clarification requests from the model, thereby reducing the need for manual intervention. Final outputs are enclosed in XML-style tags (e.g., <final_response>) and are parsed into structured formats for downstream evaluation. The full FC-Agent prompt template is included in Appendix  D.1 (Listing  6). This architecture emphasizes traceability, minimal supervision, and broad applicability across various SOP domains.
A.2 ReACT Agent
Built on the LangChain ReAct framework [30], this agent is designed to handle SOPs involving complex branching logic, intermediate reasoning, and decision pathways. It ingests the SOP text, structured task input, and tool specifications JSON, similar to the FC Agent. Tools are dynamically discovered and wrapped using @tool annotation to ensure compatibility with the LangChain agent execution framework. The ReAct Agent interleaves reasoning traces (Thoughts) with explicit tool calls (Actions), which are executed via LLM’s tool-calling interface. The resulting Observations are fed back into the reasoning loop, enabling iterative planning and dynamic adaptation to SOP-specific control flows. This process continues until a conclusive Final Answer is reached, marked using tags such as <final_decision> and <final_decision_reason>. The system and user prompt templates used to steer the ReAct Agent are detailed in Appendix  D.2 (Listings  7 and  8). This agent architecture supports more sophisticated workflows, enabling adaptive behavior in SOPs requiring fallback logic, verification steps, or multi-phase reasoning.

### v2

@@ 4.1. Evaluation Methodology
We evaluate FC-Agent and ReAct-Agent on SOP-Bench using three inputs: (1) Task, (2) SOP (instructions), and (3) ToolSpecs. The outputs include (1) intermediate tool outputs, (2) final response, and (3) trace of tool use and reasoning. We define: $n$ (total number of tasks per SOP), $n_{c}$ (number of tasks per SOP marked complete by the agent), and $n_{a}$ (number of correctly completed tasks per SOP). Evaluation metrics are: (1) Execution Completion Rate ($ECR$): proportion of tasks the agent marked as complete, $ECR=\frac{n_{c}}{n}$; (2) Conditional Task Success Rate ($C$-TSRTSR): fraction of completed tasks that match the ground truth, $C$-$TSR=\frac{n_{a}}{n_{c}}$; and (3) Task Success Rate ($TSR$): overall accuracy, $TSR=\frac{n_{a}}{n}$. Though this work reports only $ECR$ and $TSR$, future research should incorporate tailored metrics to better assess steps planning, tool mapping, tool calling accuracies, to better understand agent performance bottlenecks. We welcome the research community to contribute such agent evaluation metrics to SOP-Bench.
For all evaluations, we have kept the LLM parameters temperature at 0.5 and output token limits at 8000.
5. Results
5.1. Results and Insights
We evaluate the two agents, FC and ReAct, with 11 different LLMs in our experiments: Llama-3.3-70B-Instruct (Meta, 2024), GPT-OSS-120B (OpenAI, 2025), DeepSeek-R1 (DeepSeek, 2025), Claude 3.5 Sonnet v2 (Anthropic, 2024), Claude 3.7 Sonnet (Anthropic, 2025a), Claude 4 Sonnet (Anthropic, 2025b), Claude 4 Opus (Anthropic, 2025b), Claude 4.1 Opus (Anthropic, 2025b), Claude Sonnet 4.5 (Anthropic, 2025d), and Claude Opus 4.5 (Anthropic, 2025c). These models were selected based on availability within our organization and their known strong performance on agentic tasks, explaining the focus on larger, frontier models rather than smaller alternatives. Detailed experiment results of each SOP, Agent and LLM are in Appendix B.
5.1.1. LLM–Agent Architecture Must Be Task Dependent
Our evaluation shows that no single agent architecture-model combination consistently dominates across SOP-Bench, highlighting the need for task-dependent architecture selection in SOP automation.
Using Claude 4 Opus as a controlled comparison, Table 4 contrasts Function-Calling (FC) and ReAct agents across all 13 SOPs. ReAct achieves a higher average Task Success Rate (TSR) than FC (72% vs. 68%), but at a significant latency cost: 87.82s $\pm$ 10.30s per task compared to 66.21s $\pm$ 9.57s for FC, a 33% increase. Moreover, ReAct outperforms FC on only 8 of 13 SOPs, indicating that average gains obscure strong task-level variation.
Architecture–Model Co-Design Is Non-Monotonic. Performance does not scale monotonically with stronger models, particularly for reasoning-centric architectures. FC agents show modest, consistent improvements (Claude 3.5 v2: 65.8.2% $\rightarrow$ Claude 4.5 Sonnet: 67.5% TSR), while ReAct agents degrades: Claude 3.5 v2 achieves 65.4% TSR, Claude 4 Sonnet improves to 68.7%, but Claude 4.5 Sonnet drops to 63.3% (2.1-2.1 points). These results show that newer foundation models are not automatically better for existing agent architectures, and naive upgrades can silently reduce task success.
Task Characteristics Dominate Aggregate Metrics. Aggregate averages mask clear task-specific reversals. Across SOPs, FC slightly outperforms ReAct on average (65.9% vs. 61.1% TSR), yet individual tasks exhibit distinct architectural preferences. Dangerous Goods classification shows near parity (FC: 69.1%, ReAct: 69.8%), while Customer Service reveals identical failure modes (FC: 47.8%, ReAct: 47.7%). Other SOPs strongly favor one architecture, with ReAct excelling in multi-step inspection workflows and FC performing better in deterministic, tool-driven procedures.
Overall, these findings demonstrate that agent architecture must be selected based on SOP structure, tool interaction patterns, and latency constraints rather than global averages. SOP-Bench enables systematic evaluation of these trade-offs, supporting principled architecture selection for real-world procedural workflows.
Table 5. Top two performing model-agent combinations per SOP, ranked by Task Success Rate (TSR).
SOP Topic Best Second Best
Model Agt TSR Model Agt TSR
Patient Intake Claude 4.1 Opus ReAct 1.00 Claude 4.1 Opus FC 1.00
Aircraft Inspection Claude 3.7 Sonnet ReAct 0.99 Claude 4.5 Opus ReAct 0.97
Email Intent Claude 4 Sonnet ReAct 0.99 Claude 4.1 Opus FC 0.98
Referral Abuse Easy Claude 3.5 v2 Sonnet ReAct 0.98 Claude 3.5 v2 Sonnet FC 0.97
Referral Abuse Hard Claude 4 Opus FC 0.98 Claude 4.5 Opus FC 0.97
Video Classification Claude 4 Sonnet FC 0.95 Claude 4 Sonnet ReAct 0.95
Dangerous Goods Claude 4 Sonnet FC 0.87 Claude 4.1 Opus ReAct 0.83
Traffic Spoofing Claude 4.5 Sonnet FC 0.86 Claude 4 Sonnet FC 0.79
Customer Service Llama 3.3 70B ReAct 0.79 Claude 4.5 Opus FC 0.71
Warehouse Inspection Claude 4.1 Opus ReAct 0.69 Claude 4 Opus ReAct 0.65
Content Flagging Deepseek R1 ReAct 0.60 Claude 3.5 v2 Sonnet ReAct 0.60
Video Annotation GPT OSS 120B ReAct 0.58 Deepseek R1 ReAct 0.58
Know Your Business Claude 4.5 Opus ReAct 0.58 Claude 4.5 Haiku FC 0.57
5.1.2. Domain Variance Exposes Generalization Limits
Agent performance on SOP-Bench varies substantially across procedural domains, exposing clear limits in cross-domain generalization. Table 5 shows the top two model–agent combinations per SOP and confirms that no single configuration dominates across tasks. Best-case performance ranges from near-perfect accuracy on Patient Intake (100% TSR) to only 57% on Know Your Business, underscoring the necessity of domain-specific evaluation. Aggregated results in Table 6 further illustrate this variability: the FC agent achieves its highest average performance with Claude 4 Opus (68.2% TSR), while ReAct performs best with Claude 4 Opus as well (73.8% TSR). These rankings reflect a fixed agent configuration per model; further agent–model co-optimization could alter absolute performance but does not change the observed domain sensitivity.
[…]
Harder SOPs also exhibit greater performance variance across models. Video Annotation shows the highest standard deviation in TSR (13.8%), indicating brittle and inconsistent agent behavior in long-horizon, tool-intensive workflows. Overall, these results demonstrate that success on one domain does not reliably transfer to others, validating SOP-Bench’s design goal of capturing orthogonal dimensions of procedural reasoning, tool use, and constraint adherence.
Table 6. Average Task Success Rate (TSR) by model and agent type, averaged across all SOPs. FC agent was only evaluated with Claude models due to its design for native tool calling. ReAct agent was evaluated across all models to compare reasoning capabilities.
ReAct Agent FC Agent
Model (rank) TSR Model (rank) TSR
Claude 4 Opus (1) 0.724 Claude 4 Opus (1) 0.682
Claude 4.1 Opus (2) 0.694 Claude 4 Sonnet (2) 0.678
Claude 4 Sonnet (3) 0.687 Claude 4.5 Sonnet (3) 0.675
Claude 4.5 Opus (4) 0.683 Claude 4.5 Opus (4) 0.667
Claude 3.5 v2 Sonnet (5) 0.654 Claude 4.1 Opus (5) 0.664
Claude 4.5 Sonnet (6) 0.633 Claude 3.5 v2 Sonnet (6) 0.658
Llama 3.3 70B (7) 0.565 Claude 4.5 Haiku (7) 0.624
Deepseek R1 (8) 0.557 Claude 3.7 Sonnet (8) 0.623
Claude 4.5 Haiku (9) 0.522
Claude 3.7 Sonnet (10) 0.518
GPT OSS 120B (11) 0.484
5.2. Ablation studies
5.2.1. Tool selection under registry bloat
To investigate the impact of tool registry size on agent performance, we conducted a controlled ablation study on the Video Annotation SOP using Claude Sonnet 4.5. We evaluated agent(react) performance under two conditions: (1) Bloated Registry: 26 tools (6 task-relevant + 20 distractor tools), and (2) Minimal Registry: 6 tools (task-relevant only). The minimal registry condition achieved 37% TSR compared to 20.8% TSR in the bloated condition, representing a 1.7x improvement. This suggests tool registry bloat could be a performance bottleneck: introducing 20 distractor tools reduced success rates by 16.2 percentage points despite agents having access to all necessary tools. Task-specific tool pruning may be necessary for practical deployment.
5.2.2. Controlled Complexity Variation in Referral Abuse Detection
We conduct a controlled ablation on the Referral Abuse Detection SOP by comparing its Easy and Hard variants, which share identical task objectives, inputs, and tool interfaces but differ in logical complexity (Appendix B). As shown in Table 4, Claude 4 Opus maintains similarly high TSR on both variants for both FC and ReAct agents, indicating robustness to increased procedural complexity. However, this accuracy comes at a clear reasoning cost: execution time increases by 51.5% for the FC agent (53.99s $\rightarrow$ 81.79s) and by 34.2% for the ReAct agent (66.45s $\rightarrow$ 89.16s) on the Hard SOP.
5.2.3. Long-Context Extraction, Not Conditional Reasoning, could limit Agent Performance
We observe variability in how context length and branching complexity affect performance. Long-context SOPs (e.g., Video Annotation: 4,492 tokens, 8 sequential steps, 2 branching decision points) show lower TSR (0.28), while high-branching SOPs (e.g., Referral Abuse Detection Hard: 2576 tokens, 28 sequential steps, 12 decision points) maintain high TSR (0.84). However, these SOPs differ in multiple dimensions (tool count, domain), preventing definitive conclusions. Further investigation with controlled SOP variants is needed to isolate the independent effects of context length versus branching complexity.
6. Conclusions and Next Steps
We present SOP-Bench, a comprehensive benchmark of 2,000+ executable tasks derived from human expert-authored SOPs across 12 industrial domains. SOP-Bench enables systematic evaluation of LLM agents under realistic procedural constraints, supporting principled investigation of agent architectures, model capabilities, and deployment strategies. Our baseline experiments across representative frontier models validate a central finding: agent performance is highly task- and context-dependent, making rigorous evaluation on realistic SOPs essential prior to production deployment.
Our results show that newer models do not necessarily yield better outcomes (e.g., Claude 4 family: 73.8% vs. Claude 4.5 family: 58.8% average TSR on ReAct), and that no single model–agent combination dominates across domains (TSR ranging from 57% to 100%). These findings illustrate how SOP-Bench can be used to analyze architecture–model interactions, surface performance bottlenecks, and inform deployment decisions without costly production experimentation. Rather than serving as a leaderboard, SOP-Bench provides an evaluation substrate for targeted scientific inquiry. It enables researchers to study procedural ambiguity, tool selection, and complexity scaling, while allowing practitioners to validate architecture choices against domain-specific requirements before deployment.
Looking ahead, we plan to extend SOP-Bench along several dimensions, while actively inviting community participation. We will introduce controlled variants of the same SOP with increasing complexity, incorporate multimodal instructions such as images and tables, and model hierarchical SOPs with nested structure and context switching. Beyond these extensions, we invite the community to leverage our human–AI collaborative methodology to author new SOPs, generate executable artifacts, and contribute domain-specific tasks back to SOP-Bench. This enables the benchmark to evolve organically, expand to new industries, and serve as a shared evaluation substrate for studying agent behavior.
We release the complete benchmark, baseline agents, and evaluation framework at https://github.com/amazon-science/sop-bench.
@@ 6. Conclusions and Next Steps
Use of Generative AI
This work employed generative AI models as part of a human-AI collaborative framework for benchmark construction. Specifically, we used Anthropic’s Claude 3.5 Sonnet v2 for the following purposes:
Dataset Schema Generation: Claude 3.5 Sonnet v2 generated structured dataset schemas from human-authored SOPs and task contexts using one-shot prompting. These schemas specified input parameters, decision points, and expected outcomes.
SOP Refinement: The model refined human-authored SOPs for consistency and completeness, ensuring alignment with dataset schemas while preserving domain-specific procedural logic.
Artifact Generation: Claude 3.5 Sonnet v2 generated supporting artifacts including: (1) synthetic datasets representing full procedural contexts with diverse decision pathways and edge cases, (2) API specifications defining tool inputs and outputs, (3) tool specifications for agent interaction, and (4) executable tool code implementing procedural logic.
Complexity Assessment: Claude 3.5 v2 provided complexity ratings for each SOP on a 1-10 scale, calibrated against Berkeley function calling tasks as baseline.
Human Validation: All AI-generated content underwent rigorous human validation by domain experts. Human experts: (1) authored all original SOPs based on real-world industrial procedures, (2) validated dataset schemas for completeness and domain accuracy, (3) reviewed and corrected SOP refinements for procedural correctness, (4) validated datasets for logical consistency and appropriate data types, (5) verified API and tool specifications aligned with dataset schemas, (6) executed and validated tool code for correctness, and (7) introduced controlled complexity to mirror real-world conditions. The human-AI collaborative approach ensured procedural authenticity while maintaining reproducibility and scalability.
Evaluation: The benchmark evaluation itself used 11 different LLMs (Claude 3.5 v2 Sonnet, Claude 3.7 Sonnet, Claude 4 Sonnet, Claude 4 Opus, Claude 4.1 Opus, Claude 4.5 Haiku, Claude 4.5 Sonnet, Claude 4.5 Opus, DeepSeek R1, GPT OSS 120B, Llama 3.3 70B) as the agents being evaluated, not as tools for analysis or writing.
Manuscript Writing: Claude 4.5 Sonnet was used to correct grammatical errors in the manuscript. All technical content, experimental design, analysis, and conclusions were authored by the human researchers.
Main Image: The main figure of the paper, which presents an overview of the Human–AI Collaborative workflow for SOP-Bench, was initially generated using ChatGPT, with the paper content provided as grounding material. Since the automatically generated figure did not fully capture the required structural and cosmetic refinements, the authors subsequently refined the figure manually to accurately represent the workflow, clearly delineate the Human and AI steps, and ensure faithful depiction of the overall process.
@@ Appendix A Technical Details on Agents
A.1. FC Agent (Function-Calling Agent)
This is a prompt-driven, lightweight execution engine leveraging the native tool calling capability of LLMs (e.g., Claude via AWS Bedrock). It accepts an SOP text and a structured task input, dynamically generates prompts using a parameterized template engine, and maintains a conversation loop with the LLM throughout execution. During agent initialization, it registers external tools by converting them from OpenAPI-style tool specifications into Bedrock’s function-calling format. During execution, when the model triggers a tool_use event, the agent extracts the relevant tool name and parameters, executes the corresponding function via a domain-specific ToolManager, and delivers the output to the model through a tool_result message. The agent implements an iterative conversation pattern with a maximum of 10 iterations, where each iteration processes tool calls sequentially until no further tool uses are detected. Final outputs are enclosed in XML-style tags (e.g., ¡final_decision¿) and are parsed into structured formats for downstream evaluation. This architecture emphasizes traceability, minimal supervision, and broad applicability across various SOP domains. Note that this agent is incompatible with Amazon Nova models, which do not support native function calling.
A.2. ReAct Agent
Built on a custom implementation of the ReAct framework [Yao et al., 2023], this agent is designed to handle SOPs involving complex branching logic, intermediate reasoning, and decision pathways. It ingests the SOP text, structured task input, and tool specifications, similar to the FC Agent. Tools are dynamically discovered from the ToolManager and their descriptions are formatted with detailed parameter information including types, requirements, and allowed values. The ReAct Agent interleaves reasoning traces (Thoughts) with explicit tool calls (Actions), which are executed via the ToolManager. The resulting Observations are fed back into the reasoning loop, enabling iterative planning and dynamic adaptation to SOP-specific control flows. The agent implements strict enforcement mechanisms to prevent premature final answers without tool usage, ensuring that at least one tool is executed before providing conclusions. Response parsing uses regular expressions to extract Thought, Action, Action Input, and Final Answer components, with robust error handling for malformed JSON inputs. This process continues for up to 15 iterations until a conclusive Final Answer is reached, marked using tags such as ¡final_decision¿. The agent maintains comprehensive execution traces including intermediate steps, tool call records with success status, and iteration counts. This agent architecture supports sophisticated workflows, enabling adaptive behavior in SOPs requiring fallback logic, verification steps, or multi-phase reasoning.